OpenStudy (anonymous):

NEED HELP PLEASE <3 Mary invested a certain amount of money in a savings account paying 3% interest every year. when she withdrew her money at the end of 3 years, she received 180 in interest. how much money did mary place in the savings account?

OpenStudy (anonymous):

Simple interest? Compound interest?

OpenStudy (anonymous):

simple

OpenStudy (anonymous):

So, P * 0.03 * 3 = 180 => P = 180/(3*0.03) = 180/0.09 = 18000/9 = 2000

OpenStudy (anonymous):

i still don't understand

OpenStudy (anonymous):

Well.....if she invested amount P and earned 3% (0.03 times that amount) per year for 3 years, her total interet amount will be P times 3% times 3. You are asked to find P where she earned $180 in total interest. So, P * 0.03 * 3 = 180 => P = 180/(3*0.03) => P = 180/0.09 => P = 18000/9 = 2000

OpenStudy (anonymous):

so she invested $2000

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

A medal is the best way to say thanks rather than type that up. :)

OpenStudy (anonymous):

:) kay